Please show equations not finance calc

A semi-annual coupon bond has a 6 percent coupon rate, a $1,000 face value, a current value of $1,036.09, and 3 years until the first call date. What is the call price if the yield to call is 6.5 percent?
Answer #1

YTC = Couporadalur Pure forte YTC = - Coupon + 9 (CallValue-Purchase Price) noof years (CallValue+Purchase Price)

Let Call Value be C

60+ (C-1036.09) 0.065 = (C+1036.09)

C + 1036.09 180 + C - 1036.09) 0.065*

0.09750 + 101.018775 = 0 - 856.09

0.9025 C = 957.108775

C = 957.108775/ 0.9025

= $ 1060.51

Call Value is $ 1060.51

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.